Question 6

Cardiff and Delp is an architectural firm that provides services for residential construction projects. The following data peOverhead Assigned
Answer #1

Note---

rate =A/B = expect cost/expect activity driver

Overhead assigned= D*C = activity driver incurred* activity rate.(A) overlead assig uponerende $ 850do $ 18810 5oo howed houl (8) (c) Activity Espected costs espected activing Setivity Activ